A ball is thrown upward from ground level. Its height, h, in feet above the ground after t seconds os h=48t-16t^2. Find the maximum height of the ball. Find the time it takes for the ball to fall back to the ground.

Answer by nerdybill(7384)

A ball is thrown upward from ground level. Its height, h, in feet above the ground after t seconds os h=48t-16t^2. Find the maximum height of the ball. Find the time it takes for the ball to fall back to the ground.
.
Max height is at the vertex.
Value of t at the vertex is found by:
t = -b/(2a)
t = -48/(2(-16))
t = -48/(-32)
t = -48/(-32)
t = 1.5 seconds
height at that time is:
h=48(1.5)-16(1.5)^2
h=72-36
h=36 feet (answer to first part)
.
Time to fall back to ground -- set h to 0 and solve for t:
h=48t-16t^2
0=48t-16t^2
0=3t-t^2
0=t(3-t)
t = {0, 3}
throw out the zero leaving:
t = 3 seconds (answer to second part)